# Semi-empirical Modeling of Supermassive Black Hole Evolution: Highlighting a possible tension between Demographics and Gravitational Wave Background

### v1

@@ 3.1.3 GW stochastic background
In principle, one could envisage physically-motivated modifications of the current setup to potentially ease the tension between the PTA signal and the model predictions. For example, a local BH mass function more skewed toward larger masses than the ones presented in this work could provide room for more frequent and/or more massive BH mergers, thus boosting the implied GW background signal. This could originate from a particularly flat stellar mass function at the high-mass end, and/or from an appreciably high normalization of the $M_{\bullet}-\sigma$ relationship (as may be the case for cored ellipticals, see [196]; also K. M. Varadarajan in preparation). Another possibility may involve a radiative efficiency increasing with the BH mass (see [197, 23]) which progressively reduces the number density of the most massive BHs, so as to reconcile the local BH mass function with the one derived from accretion even in the event of an increased number of BH mergers. However, a couple of reasons suggest that these attempts are likely doomed to fail. First, the merger rate density of galaxies is nowadays well constrained by JWST data, and even requiring the merger rate density of BHs to mirror it is not sufficient to reproduce the PTA background, as shown by our analysis. Therefore, independently of the shape of the local BH mass function, explaining all the GW background via BH mergers would require to assume that the merger rate per BH is appreciably larger than that of galaxies. Second, leaving more room to mergers at late times in building up the local BH mass function at the high-mass end could be dangerous, since mergers would also tend to reduce the number of low-mass BH (which merge into larger ones) so distorting the mass function there. All in all, solving the tension by acting solely on the accretion-related parameters would require some fine-tuning and would remain a very challenging task. We plan to perform an in-depth analysis of these and further possibilities in a dedicated future work.
@@ 4 Summary
  • Different local mass function determinations imply substantially different values for the parameters regulating the accretion mode of the supermassive BH population. The BH mass function based on the stellar mass function and observed $M_{\bullet}-M_{\star}$ relationship requires low radiative efficiency $\epsilon\sim 0.06$, small average values of the Eddington ratio $\lambda\sim 10^{-2}$ with a rather strong dependence on redshift $\lambda\propto(1+z)^{2.5}$; a high dispersion $\sigma_{\log M_{\rm H}}\sim 0.7$ in the relationship between BH and host halo mass is also required to reproduce the BH clustering. The BH mass function based on the velocity dispersion function and the observed $M_{\bullet}-\sigma$ relationship allows for standard values of the efficiency $\epsilon\sim 0.1$, local average Eddington ratios $\lambda\lesssim 10^{-1}$ with a mild redshift evolution $\lambda\propto(1+z)$, and scatter of $\sigma_{\log M_{\rm H}}\sim 0.6$ in the BH-halo mass relation. Finally, the BH mass determination based on the debiased $M_{\bullet}-\sigma$ relationship requires considerably high values of the efficiency $\epsilon\sim 0.3$, local average Eddington ratio $\lambda\lesssim 0.3$ modestly evolving with redshift $\lambda\propto(1+z)^{0.4}$, and appreciably lower scatter $\sigma_{\log M_{\rm H}}\sim 0.3$ in the BH vs. halo mass relation. The difference in efficiency could have far-reaching implications on the BH spinning state of the BH population, though systematic uncertainties in the local BH mass function determination do not allow to draw firm conclusions yet.
  • In all the above instances the fits from our semi-empirical model to the local BH mass function and BH clustering are remarkably good. Moreover, the outcomes of our analysis are able to reproduce auxiliary observables that were not included in the fitting procedure: the active BH mass function at high redshift $z\sim 3-6$ comprising the recently discovered population of little red dots, the local BH vs. halo mass relationship, the local BH duty cycle and the evolution of the mean Eddington ratio across cosmic times. Such an agreement supports our semi-empirical, data-driven model and the outputs of the continuity plus Smoluchowski equation approach.
@@
In a future perspective, it would be beneficial to input our semi-empirical model with improved determinations of the AGN luminosity functions and active BH mass functions, especially toward high redshift. In this respects, new data from Euclid, JWST, e-ROSITA and eventually NewATHENA will clear residual uncertainties on bolometric corrections, obscured AGN fraction, and abundance of BHs with faint bolometric luminosities. In addition, it would be extremely important to solve the longstanding debate concerning the local BH mass function determination. Although to build a BH mass function based on an extended sample of quiescent BHs with robustly determined masses would remain challenging in the near future, a great step forward could be to identify the most fundamental BH-galaxy relationship and assess the observational and systematics issues that may affect and bias it. Wide galaxy surveys that will be conducted by the LSST and eventually by ELT will help in this respect. All the above will improve appreciably the constraints on the parameters regulating the accretion mode for the growth of the BH population, like the AGN luminosity function, the radiative efficiency (hence BH spins), and the redshift-dependent Eddington ratio distributions.
As for the impact of binary BH mergers on the growth of the supermassive BH population, in the future it will be effectively constrained by more precise measurements of the nano-Hz GW background from PTA experiments; this will be possible thanks to the advent of the SKAO and its precursors, that will allow the monitoring of a substantially increased amounts of pulsars. These data will greatly improve the constraints on the mass and possibly the redshift dependence of the BH merger rate, allowing to confirm the conclusion of the present analysis concerning a modest contribution of binary BH mergers to the GW background. If so, the implications would be enormous. On the one hand, this would be a good news for LISA, since supermassive BH mergers may originate an appreciable background noise in the mHz band, degrading the effective sensitivity of the antennae to many primary science goals of the mission (e.g., [198]), like the detection of individual massive BH binaries, verified galactic binaries, and extreme mass ratio inspirals. On the other hand, our current results could motivate an exciting search for alternative astrophysical sources or even rise the hype for a primordial, cosmological origin of the remaining signal measured by PTA experiments.
@@ Acknowledgments
This work was partially funded from the projects: INAF GO-GTO Normal 2023 funding scheme with the project "Serendipitous H-ATLAS-fields Observations of Radio Extragalactic Sources (SHORES)"; INAF Large Grant 2022 funding scheme with the project "MeerKAT and LOFAR Team up: a Unique Radio Window on Galaxy/AGN co-Evolution; “Data Science methods for MultiMessenger Astrophysics & Multi-Survey Cosmology” funded by the Italian Ministry of University and Research, Programmazione triennale 2021/2023 (DM n.2503 dd. 9 December 2019), Programma Congiunto Scuole; Italian Research Center on High Performance Computing Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2 within the activities of Spoke 3 (Astrophysics and Cosmos Observations); European Union — NextGenerationEU under the PRIN MUR 2022 project n. 20224JR28W “Charting unexplored avenues in Dark Matter”. MB acknowledges that this article was produced while attending the PhD program in PhD in Space Science and Technology at the University of Trento, Cycle XXXIX, with the support of a scholarship financed by the Ministerial Decree no. 118 of 2nd March 2023, based on the NRRP - funded by the European Union - NextGenerationEU - Mission 4 "Education and Research", Component 1 "Enhancement of the offer of educational services: from nurseries to universities” - Investment 4.1 “Extension of the number of research doctorates and innovative doctorates for public administration and cultural heritage” - CUP E66E23000110001. LB acknowledges financial support from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) STRUCTURES. HF acknowledges support at Fudan University from the Shanghai Super Post-doctoral Excellence Program grant No. 2024008. DR acknowledges support from the University of Southampton via the Mayflower studentship.
Appendix A Details on the input AGN luminosity function
In this Appendix we report some details on the AGN bolometric luminosity function employed in our semi-empirical model. Fiducially, we use the analytic rendition to the multi-wavelength data in optical, UV, IR and X-ray bands by [74], in terms of the fit labeled ’A’ in their Table 4. This is expressed as
$N(\log L_{\rm AGN})=\frac{\phi_{\star}}{(L_{\rm AGN}/L_{\star})^{\gamma_{1}}+(L_{\rm AGN}/L_{\star})^{\gamma_{2}}}$ (A.1)
in terms of the parameter set $\{\phi_{\star}(z),\gamma_{1}(z),\gamma_{2}(z),L_{\star}(z)\}$. The latter depend on redshift $z$ according to the expressions
$\log\phi_{\star}\,[{\rm Mpc}^{-3}\,{\rm dex}^{-1}]=d_{0}+d_{1}\,(1+z)$ (A.2)
$\gamma_{1}=a_{0}+a_{1}\,(1+z)+a_{2}\,[2\,(1+z)^{2}-1]$
$\gamma_{2}=\frac{2\,b_{0}}{[(1+z)/3]^{b_{1}}+[(1+z)/3]^{b_{2}}}$
$\log L_{\star}\,[L_{\odot}]=\frac{2\,c_{0}}{[(1+z)/3]^{c_{1}}+[(1+z)/3]^{c_{2}}}$
where $\{a_{0},a_{1},a_{2}\}=\{0.8569,-0.2614,0.02\}$, $\{b_{0},b_{1},b_{2}\}=\{2.5375,-1.0425,1.1201\}$, $\{c_{0},c_{1},c_{2}\}=\{13.0088,-0.5759,0.4554\}$, $\{d_{0},d_{1}\}=\{-3.5426,-0.3936\}$. The resulting bolometric AGN luminosity functions are plotted as a function of redshift over the range $z\sim 0-7$ (color-coded) in the top panel of Figure 11.
@@ Appendix A Details on the input AGN luminosity function
Admittedly, the above fitting formula Equation (A.1) by [74] currently constitutes the best homogenized rendition of different multi-wavelength datasets on the AGN luminosity function, converted to bolometric and accounting for obscured fraction with common corrections. However, recently the shape of the [74] luminosity function at $z\gtrsim 4.5$ has been somewhat challenged by new estimates from optical and X-ray data (see [199, 183, 184, 190, 191]). The issue has been also reinforced by the population of overmassive, X-ray underluminous BHs recently discovered by JWST (e.g., [82, 83, 84, 85, 86, 200]), partly located in the so called ‘little red dots’ sources (see [87, 88, 89, 90, 91, 92]).
@@
The current situation may be visualized in the lower panel of Figure 11, where the AGN luminosity function by [74] at redshift $z\sim 5-7$ (color-coded) is illustrated and compared to the most recent data collected by [191]. Besides the still large uncertainties in the data, there is a tendency for the fit by [74] to somewhat under-predict the observational estimates. Remarkably, we checked that the rendition by [74] can be reconciled with these new data just by changing the parameter controlling the number density evolution $d_{1}\rightarrow d_{1}+0.2$ for $z\gtrsim 4.5$ (i.e., the negative number density evolution is slowed down). The results are illustrated by the solid lines, which agree with the data remarkably better.
@@
One may wonder what would be the impact of the revised fit to the AGN luminosity function on the results presented in the main text. We have checked that this is irrelevant for the building up of the local BH mass function, since most of the accreted mass is acquired at redshift $z\lesssim 4$. In addition, the correction also turns out to be negligibly important for the GW background in the nano-Hz band, since most of it is originated at $z\lesssim 3$ (see Fig. 9). A more quantitative analysis is presented in Appendix B.
@@ Appendix B Additional tests
In this Appendix we present additional tests on the robustness of the result on the expected GW background to variations in some assumptions of our analysis. We focus on the local BH mass function determination based on the stellar mass function and the $M_{\bullet}-M_{\star}$ relation, since this tends to maximize the impact of binary BH mergers. We show how the posterior on the GW normalized stretch amplitude is altered when the following assumptions are modified with respect to our fiducial analysis in the main text.
[…]
  • Kinetic efficiency $\epsilon_{\rm kin}$. In the main text we have assumed $\epsilon_{\rm kin}=0.15\,\epsilon_{\rm thin}$ following the analysis by [135]; given the bestfit values of $\epsilon_{\rm thin}\lesssim 0.1$, rather small kinetic efficiencies are implied. However, since it has been claimed (e.g., [136, 137]) that the kinetic efficiency could be larger, here we test the impact of assuming $\epsilon_{\rm kin}=0.2$.

### v2

@@ 3.1.3 Stochastic GW background
In principle, one could envisage physically-motivated modifications of the current setup to potentially ease the tension between the PTA signal and the model predictions. For example, a local BH mass function more skewed toward larger masses than the ones presented in this work could provide room for more frequent and/or more massive BH mergers, thus boosting the implied GW background signal. This could originate from a particularly flat stellar mass function at the high-mass end, and/or from an appreciably high normalization of the $M_{\bullet}-\sigma$ relationship (as may be the case for cored ellipticals, see [Saglia2024]; also K. M. Varadarajan in preparation). Another possibility may involve a radiative efficiency increasing with the BH mass (see [Cao2008, Shankar2013]) which progressively reduces the number density of the most massive BHs, so as to reconcile the local BH mass function with the one derived from accretion even in the event of an increased number of BH mergers. However, a couple of reasons suggest that these attempts are likely doomed to fail. First, the merger rate density of galaxies is nowadays well constrained by JWST data, and even requiring the merger rate density of BHs to mirror it is not sufficient to reproduce the PTA background, as shown by our analysis. Therefore, independently of the shape of the local BH mass function, explaining all the GW background via BH mergers would require to assume that the merger rate per BH is appreciably larger than that of galaxies. Second, leaving more room to mergers at late times in building up the local BH mass function at the high-mass end could be dangerous, since mergers would also tend to reduce the number of low-mass BH (which merge into larger ones) so distorting the mass function there. All in all, solving the tension by acting solely on the accretion-related parameters would require some fine-tuning and would remain a very challenging task. We plan to perform an in-depth analysis of these and further possibilities in a dedicated future work.
@@ 4 Summary
  • Different local mass function determinations imply substantially different values for the parameters regulating the accretion mode of the supermassive BH population. The BH mass function based on the stellar mass function and observed $M_{\bullet}-M_{\star}$ relationship requires low radiative efficiency $\epsilon\sim 0.06$, small local average values of the Eddington ratio $\lambda\sim 10^{-2}$ with a rather strong dependence on redshift $\lambda\propto(1+z)^{2.5}$; a high dispersion $\sigma_{\log M_{\rm H}}\sim 0.7$ in the relationship between BH and host halo mass is also required to reproduce the BH clustering. The BH mass function based on the velocity dispersion function and the observed $M_{\bullet}-\sigma$ relationship allows for standard values of the efficiency $\epsilon\sim 0.1$, local average Eddington ratios $\lambda\lesssim 10^{-1}$ with a mild redshift evolution $\lambda\propto(1+z)$, and scatter of $\sigma_{\log M_{\rm H}}\sim 0.6$ in the BH-halo mass relation. Finally, the BH mass determination based on the debiased $M_{\bullet}-\sigma$ relationship requires considerably high values of the efficiency $\epsilon\sim 0.3$, local average Eddington ratio $\lambda\lesssim 0.3$ modestly evolving with redshift $\lambda\propto(1+z)^{0.4}$, and appreciably lower scatter $\sigma_{\log M_{\rm H}}\sim 0.3$ in the BH vs. halo mass relation. The difference in efficiency could have far-reaching implications on the BH spinning state of the BH population, though systematic uncertainties in the local BH mass function determination do not allow to draw firm conclusions yet.
  • In all the above instances the fits from our semi-empirical framework to the local BH mass function and BH clustering are remarkably good. Moreover, the outcomes of our analysis are able to reproduce auxiliary observables that were not included in the fitting procedure: the active BH mass function at high redshift $z\sim 3-6$ comprising the recently discovered population of little red dots, the local BH vs. halo mass relationship, the local BH duty cycle and the evolution of the mean Eddington ratio across cosmic times. Such an agreement supports our semi-empirical, data-driven framework and the outputs of the continuity equation approach.
@@
In a future perspective, it would be beneficial to input our semi-empirical framework with improved determinations of the AGN luminosity functions and active BH mass functions, especially toward high redshift. In this respects, new data from Euclid, JWST, e-ROSITA and eventually NewATHENA will clear residual uncertainties on bolometric corrections, obscured AGN fraction, and abundance of BHs with faint bolometric luminosities. In addition, it would be extremely important to solve the longstanding debate concerning the local BH mass function determination. Although to build a BH mass function based on an extended sample of quiescent BHs with robustly determined masses would remain challenging in the near future, a great step forward could be to identify the most fundamental BH-galaxy relationship and assess the observational and systematics issues that may affect and bias it. Wide galaxy surveys that will be conducted by the Rubin-LSST and eventually by ELT will help in this respect. All the above will improve appreciably the constraints on the parameters regulating the accretion mode for the growth of the BH population, like the AGN luminosity function, the radiative efficiency (hence BH spins), and the redshift-dependent Eddington ratio distributions.
As for the impact of binary BH mergers on the growth of the supermassive BH population, in the future it will be effectively constrained by more precise measurements of the nano-Hz GW background from PTA experiments; this will be possible thanks to the advent of the SKAO and its precursors, that will allow the monitoring of a substantially increased amounts of pulsars. These data will greatly improve the constraints on the mass and possibly the redshift dependence of the BH merger rate, allowing to confirm the conclusion of the present analysis concerning a modest contribution of binary BH mergers to the GW background. If so, the implications would be enormous. On the one hand, this would be a good news for LISA, since supermassive BH mergers may originate an appreciable background noise in the mHz band, degrading the effective sensitivity of the antennae to many primary science goals of the mission (e.g., [Huang2024]), like the detection of individual massive BH binaries, verified galactic binaries, and extreme mass ratio inspirals. On the other hand, our current results could motivate an exciting search for alternative astrophysical sources or even rise the hype for a primordial, cosmological origin of the remaining signal measured by PTA experiments.
@@ Acknowledgments
We thank the anonymous referee for constructive and helpful comments. This work was partially funded from the projects: INAF GO-GTO Normal 2023 funding scheme with the project "Serendipitous H-ATLAS-fields Observations of Radio Extragalactic Sources (SHORES)"; INAF Large Grant 2022 project "MeerKAT and LOFAR Team up: a Unique Radio Window on Galaxy/AGN co-Evolution; INAF Large GO 2024 project "MeerKAT and Euclid Team up: Exploring the galaxy-halo connection at cosmic noon"; “Data Science methods for MultiMessenger Astrophysics & Multi-Survey Cosmology” funded by the Italian Ministry of University and Research, Programmazione triennale 2021/2023 (DM n.2503 dd. 9 December 2019), Programma Congiunto Scuole; Italian Research Center on High Performance Computing Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2 within the activities of Spoke 3 (Astrophysics and Cosmos Observations); European Union — NextGenerationEU under the PRIN MUR 2022 project n. 20224JR28W “Charting unexplored avenues in Dark Matter”. MB acknowledges that this article was produced while attending the PhD program in PhD in Space Science and Technology at the University of Trento, Cycle XXXIX, with the support of a scholarship financed by the Ministerial Decree no. 118 of 2nd March 2023, based on the NRRP - funded by the European Union - NextGenerationEU - Mission 4 "Education and Research", Component 1 "Enhancement of the offer of educational services: from nurseries to universities” - Investment 4.1 “Extension of the number of research doctorates and innovative doctorates for public administration and cultural heritage” - CUP E66E23000110001. LB acknowledges financial support from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) STRUCTURES. HF acknowledges support at Fudan University from the Shanghai Super Post-doctoral Excellence Program grant No. 2024008. DR acknowledges support from the University of Southampton via the Mayflower studentship.
Appendix A The high-redshift AGN luminosity function
@@
In the main text we have adopted as our fiducial input the AGN luminosity function by [Shen2020], and in particular their fitting formula reported in Equation (2.2). This currently constitutes the best homogenized rendition of different multi-wavelength datasets on the AGN luminosity function, converted to bolometric and accounting for obscured fraction with common corrections. However, recently the shape of the [Shen2020] luminosity function at $z\gtrsim 4.5$ has been somewhat challenged by new estimates from optical and X-ray data (see [Wolf2021, Greene2024, Maiolino2024, Barlow2023, Barlow2025]). The issue has been also reinforced by the population of overmassive, X-ray underluminous BHs recently discovered by JWST (e.g., [Ubler2023, Furtak2023, Harikane2023, Kokorev2023, Stone2024, Maiolino2024jades]), partly located in the so called ‘little red dots’ sources (see [Kocevski2023, Matthee2024, He2024, Lai2024, Kokorev2024, Akins2025]).
@@
The current situation may be visualized in Figure 13, where the AGN luminosity function by [Shen2020] at redshift $z\sim 5-7$ (color-coded) is illustrated and compared to the most recent data collected by [Barlow2025]. Besides the still large uncertainties in the data, there is a tendency for the fit by [Shen2020] to somewhat under-predict the observational estimates. Remarkably, we checked that the rendition by [Shen2020] can be reconciled with these new data just by modifying the parameter controlling the number density evolution $d_{1}\rightarrow d_{1}+0.2$ for $z\gtrsim 4.5$ (i.e., the negative number density evolution is slowed down). The results are illustrated by the solid lines, which agree with the data remarkably better.
@@
One may wonder what would be the impact of this revised fit to the AGN luminosity function on the results presented in the main text. We have checked that this is irrelevant for the building up of the local BH mass function, since most of the accreted mass is acquired at redshift $z\lesssim 4$. In addition, the correction also turns out to be negligibly important for the GW background in the nano-Hz band, since most of it is originated at $z\lesssim 3$ (see Fig. 11). A more quantitative analysis is presented in Appendix B.
@@ Appendix B Additional tests
In this Appendix we present additional tests on the robustness of the result on the expected GW background to variations in some assumptions of our analysis. We rely on the local BH mass function determination based on the stellar mass function and the $M_{\bullet}-M_{\star}$ relation, since this tends to maximize the impact of binary BH mergers. We focus on how the posterior on the GW normalized stretch amplitude is altered when the following assumptions are modified with respect to our fiducial analysis in the main text.
  • AGN luminosity functions at high redshift. In the main text we have used the analytical rendition by [Shen2020] over the redshift range $z\sim 0-7$. Here we try to employ the correction to the $z\gtrsim 4.5$ behavior discussed in Appendix A which is suggested by the most recent highz-z data.
  • Kinetic efficiency $\epsilon_{\rm kin}$. In the main text we have assumed $\epsilon_{\rm kin}=0.15\,\epsilon_{\rm thin}$ following the analysis by [Shankar2008]; given the bestfit values of $\epsilon_{\rm thin}\lesssim 0.1$, rather small kinetic efficiencies are implied. However, since it has been claimed (e.g., [Ghisellini2013, Zubovas2018]) that the kinetic efficiency could be larger, here we test the impact of assuming $\epsilon_{\rm kin}=0.2$.
@@
Figure 1: Schematics showing the basic methodology underlying our semi-empirical framework (see Section 2 for details).
Figure 2: The AGN bolometric luminosity functions derived from multi-wavelength data by [Shen2020] (we use the fit labeled ’A’ in their Table 4; see also Section 2.2) as a function of redshift (color-coded), which is the basic input in our semi-empirical framework.
Figure 3: MCMC posterior distributions for the parameters exploited in our semi-empirical framework: radiative efficiency of the thin disk phase $\epsilon_{\rm thin}$; mean Eddington ratio $\langle\log\lambda_{0}\rangle$ and its redshift evolution parameter $\xi$; mass-dependence parameter $\eta$ in the supermassive BH merger rate and rescaling parameter $f_{\bullet\bullet\rightarrow\bullet}$ of the BH to galaxy merger rate; scatter $\sigma_{\log M_{\rm H}}$ in the abundance matching relationship between halo and BH mass. Coloured contours/lines refer to the analysis with different determinations of the local BH mass function: orange is for that obtained combining the local SMF with the $M_{\bullet}-M_{\star}$ relation, and magenta adds a narrow prior around $f_{\bullet\bullet\rightarrow\bullet}\approx 1$; blue is for that obtained combining the VDF with the observed $M_{\bullet}-\sigma$ relation; green is for that obtained combining the VDF with the debiased $M_{\bullet}-\sigma$ relation. The marginalized distributions are in arbitrary units (normalized to unity at their maximum value). Colored crosses mark bestfit values. In the uppermost panel, the gray shaded areas highlight non-physical values of the thin-disk radiative efficiency $\epsilon_{\rm thin}$ which are excluded by hard priors.
Figure 4: The supermassive BH mass function. Colored symbols refer to the local BH mass function determined from SMF and observed $M_{\bullet}-M_{\star}$ relation (green squares), from VDF and observed $M_{\bullet}-\sigma$ relation (blue circles), and from VDF and debiased $M_{\bullet}-\sigma$ relation (red rhomboids). These determinations bracket the overall systematics uncertainty on the estimates of the local BH mass function. For comparison, the dotted line is the recent determination based on the $M_{\bullet}-M_{\star}$ scaling relation by [Liepold2024]; the gray shaded area is the determination by [Shen2020], obtained convolving the AGN luminosity function with the AGN luminosity vs. Eddington ratio relationship from SDSS [Shen2009, Nobuta2012] and further correcting for obscured AGNs and inactive BHs. Grey symbols display recent estimates of the active BH mass function at $z\sim 4-6$ including BAL quasars and little red dots by [He2024] (reverse triangles), [Lai2024] (triangles), [Taylor2025] (pentagons), [Matthee2024] (hexagons), [Kokorev2024] (plus signs), and [Wu2022] (crosses). Solid lines with shaded areas illustrate the outcomes (median and $2\sigma$ credible intervals) for the local BH mass function reconstructed from our semi-empirical framework and MCMC analysis, with the same color-code as in previous Figure, while dashed lines illustrate the corresponding active BH mass functions at $z\sim 4.5$.
Figure 5: Large-scale bias as a function of BH mass. Data are from X-ray AGNs at $z\sim 0.04$ by [Powell2018, Allevato2021] (circles), from X-ray (squares) and optical (crosses) AGNs at $z\sim 0.25$ by [Krumpe2015, Krumpe2023]. Colored solid lines with shaded areas illustrate the outcomes (median and $2\sigma$ credible intervals) from our semi-empirical framework model and MCMC analysis, with the same color code as in previous Figures. The inset shows the BH bias as a function of redshift, averaged over the supermassive BH mass function. Colored lines are the outcomes from our semi-empirical framework, while symbols report observational estimates from optical (plus signs; [Shen2009, White2012, Ross2020]), X-ray (diamonds; [Allevato2011, Allevato2014, Allevato2016]) and radio-selected (inverse triangles; [Magliocchetti2017, Mazumder2022, Hale2025]) samples. Note that the redshift dependence has not been fitted upon, and constitutes just an a-posteriori validation of our semi-empirical framework.
Figure 6: Halo mass vs. BH mass relationship. Data symbols refer to estimate from clustering of X-ray (squares) and optical (circles) AGNs by [Krumpe2015], from weak lensing measurements of type-1 (inverse triangles) and type-2 (triangles) AGNs, from globular cluster dynamics (stars) by [Marasco2021], and from Tully-Fisher kinematics for reverberation-mapped AGNs by [Robinson2021] (grey shaded area). Colored solid lines with shaded areas illustrate the outcomes (median and $2\sigma$ credible intervals) from our semi-empirical framework and MCMC analysis on BH demographics and clustering, with the same color code as in previous Figures; note that the data reported here have not been fitted upon, and constitute just an a-posteriori validation. In the top left corner the colored errorbars represent the bestfit scatter $\sigma_{\log M_{\rm H}}$ on the BH to halo mass relationship inferred from our analysis.
Figure 7: Local duty cycle as a function of BH mass. Data are from AGN active fractions by [Goulding2010] (circles), [Kauffmann2003] (squares) and [Schulze2010] (diamonds). Colored solid lines with shaded areas illustrate the outcomes (median and $2\sigma$ credible intervals) from our semi-empirical framework and MCMC analysis, with the same color code as in previous Figures; note that the data reported here have not been fitted upon, and constitute just an a-posteriori validation.
Figure 8: Top panel: Average Eddington ratio vs. redshift (bottom axis) or cosmic time (top axis). Data are from: classic measurements from X-ray/mid-IR/optical AGNs by [Vestergaard2009, Nobuta2012, Dai2014, Duras2020] (circles), broad-absorption-line quasars by [He2024] (squares), [Lai2024] (triangles) [Farina2022] (diamonds) and [Fan2023] (pentagons); little red dots by [Harikane2023, Greene2024, Matthee2024, Maiolino2024] (crosses). For reference, the dashed line illustrates the mean Eddington ratio found for type-1 AGNs at $z\lesssim 4$ by [Kelly2013], as fitted by [Tucci2017]; the dotted line is the dependence assumed on phenomenological grounds by [Shankar2013]. Colored solid lines with shaded areas illustrate the outcomes (median and $2\sigma$ credible intervals) from our semi-empirical framework and MCMC analysis, with the same color code as in previous Figures; note that the data reported here have not been fitted upon, and constitute just an a-posteriori validation. The colored errorbars in the bottom left corner represent the log-normal scatter in the Eddington ratio distribution adopted in our model. Bottom panels: Eddington ratio distribution at three representative redshifts $z\approx 0.8$ (left) $1.8$ (middle) and $3.2$ (right). Circles illustrate the estimates for type-1 AGNs by [Kelly2013], plotted for the range in $\log\lambda$ where the completeness for the flux-limited SDSS sample is above 10%. Dashed line is the corresponding lognormal fit by [Tucci2017]. Colored lines with shaded areas show the outcomes of our semi-empirical framework, as above.
Figure 9: The relative contribution of binary BH mergers to the increase of the BH mass function in different mass and redshift bins (color-coded). Panels refer to the analysis with different determinations of the local BH mass function: top left is for SMF plus observed $M_{\bullet}-M_{\star}$; top right is for SMF plus observed $M_{\bullet}-M_{\star}$ plus a narrow prior centered on $f_{\bullet\bullet\rightarrow\bullet}\approx 1$; bottom left is for VDF plus observed $M_{\bullet}-\sigma$; bottom right is for VDF plus debiased $M_{\bullet}-\sigma$.
Figure 10: Normalized stretch amplitude of the gravitational wave background. Colored violins illustrate the full posterior distributions from our semi-empirical framework and MCMC analysis, with the same color code as in previous Figures (and also specified on the bottom axis); the black dots and solid lines illustrate the median and the $2\sigma$ credible interval. Grey symbols display the results from the Astrid (rhomboid; [Chen2025]) and TNG-300 numerical simulations (pentagon; [Chen2025]), and from the holodeck suite (cross; [Agazie2023]). The measurement by the NANOGrav collaboration [Agazie2023] is displayed by the yellow solid line with shaded area (median and $90\%$ credible interval).
Figure 11: The contribution to the GW stochastic background from supermassive BH mergers sliced in redshift and BH mass. Panels refer to the analysis with different determination of the local BH mass function: top left is for SMF plus observed $M_{\bullet}-M_{\star}$; top right is for SMF plus observed $M_{\bullet}-M_{\star}$ plus a narrow prior centered on $f_{\bullet\bullet\rightarrow\bullet}\approx 1$; bottom left is for VDF plus observed $M_{\bullet}-\sigma$; bottom right is for VDF plus debiased $M_{\bullet}-\sigma$. The color-scale and white contours display logarithmic values of the normalized (squared) strain parameters $\mathcal{A}^{2}$ in bins of BH mass $\log M_{\bullet}$ and redshift $z$.
Figure 12: The merger rate per descendant BH vs. redshift, referred to a reference BH mass $M_{\bullet}\sim 10^{10}\,M_{\odot}$ as a representative contributor to the GW background. Colored solid lines illustrate the outcomes from our semi-empirical framework and MCMC analysis, with the same color code as in previous Figures: the median from the posterior is reported, but the uncertainty for the red, blue and green cases essentially extends from the upper bound of the magenta line downwards (i.e., our model can only set upper limits on $f_{\bullet\bullet\rightarrow\bullet}$, see Fig. 3). The yellow lines are the reconstructed merger rate per BH from various analysis of PTA data: solid refers to [Agazie2023b], dashed to [Casey2025], and dotted to [Middleton2021]. The grey lines are illustrated for comparison: solid is the major merger rate per galaxy estimated from JWST galaxy pairs counts by [Duan2025]; dashed is the major merger rate per galaxy from the Illustris simulation by [Rodriguez2015], converted to the reference BH mass via the BH-stellar mass relation by [Kormendy2013]; dotted is the major merger rate per dark matter halo by [Fakhouri2010], converted to the reference BH mass via the BH-halo mass relation by [Ferrarese2002].
Figure 13: The AGN luminosity function at high redshift $z\gtrsim 4.5$. Colored lines refer to redshift $z\approx 5$ (blue), 66 (green), and 77 (red). Dotted lines represent the original luminosity function by [Shen2020] used as our fiducial input in the main text, while the solid lines are a modified version (see Appendix A) to account for the recent data in the range $z\sim 5-7$ from [Maiolino2024jades] (diamonds), [Greene2024] (squares), [Barlow2023] (squares), [Barlow2025] (reversed triangles) and [Wolf2021] (circle).
Figure 14: Additional tests on the contribution of binary BH mergers to the gravitational wave background. The local BH mass function determination from the stellar mass function and the $M_{\star}-M_{\bullet}$ relation is adopted. Colored violins illustrate the posterior distributions from our semi-empirical framework and MCMC analysis; the black dots and solid lines illustrate the median and the $2\sigma$ credible interval. The orange violin refers to our fiducial setting in the main text. The blue violin refers to the correction of the high-$z$ AGN luminosity function discussed in Appendix A. The green violin refers to a high value of the kinetic efficiency $\epsilon_{\rm kin}=0.2$. The magenta violin refers to the case when the PTA measurement of the GW background is not included in the likelihood. The measurement by the NANOGrav collaboration [Agazie2023] is displayed by the yellow solid line with shaded area (median and $90\%$ credible interval).